adralegab:

which expression is equivalent to sin (0.6x)-sin(0.4y)

jhonyy9:

??? wich expression - there are any choices ?

ratnakermehta20:

We know that, \[\sin C-\sin D=\cos \left(( C+D \right)/2)\sin((C-D)/2).\] \[:.\sin(0.6x)-\sin(0.4y)=\cos((0.6x+0.4y)/2)\sin((0.6x-0.4y)/2),\] \[=\cos(0.3x+0.2y)\sin(0.3x-0.2y).\]